Suppose we have a random variable X such that X = 1 with probability 1/2 and X--1 with probability 1 /2. we also have another random variable Y such that Y- X with probability 3/4 and YXwith probability 1/4. What is the covariance between them, Cov(X, Y)?

The guess 0 was marked as incorrect.
Answer #1

E(Y|X=1)=1*(3/4)+(-1)*(1/4)=1/2

E(Y|X=-1)=(-1)*(3/4)+(1)*(1/2)=-1/2

E(Y)=E(Y|X=1)*P(X=1)+E(Y|X=-1)*P(X=-1)=(1/2)*(1/2)+(-1/2)*(1/2)=0

E(X)=\sumxP(x)=1*(1/2)+(-1)*(1/2)=0

E(XY)=\sumxyP(x,y)=1*1*(1/2)*(3/4)+1*(-1)*(1/2)*(1/4)+(-1)*(-1)*(1/2)*(3/4)+(-1)*(1)*(1/2)*(1/4)=0.5

therefore Cov(X,Y)=E(XY)-E(X)*E(Y)=0.5-0*0 =0.5